OpenStudy (anonymous):

In the xy-plane, what is the slope of the line whose equation is 3x-2y= 8? A. -4 B. -8/3 C. 2/3 D. 3/2 E. 2

OpenStudy (aaronandyson):

3x-2y= 8 Compare this equation to y = mx + c: Where m = slope Therefore, 3x - 8 = 2y Therefore,3x/2 - 8/2 = y Therefore,3x/2 - 4 = y Therefore, 3/2 is the slope.

OpenStudy (aaronandyson):

Have you understood @yomamabf ?

OpenStudy (anonymous):

hang on i'm trying to process this information by trying it myself

OpenStudy (aaronandyson):

Ask doubts if any..

OpenStudy (anonymous):

Hey what is the 4?

OpenStudy (anonymous):

is that the y-intercept?

OpenStudy (aaronandyson):

Yes.

OpenStudy (anonymous):

So Y= mx+ b is Y= 3/2x-4?

OpenStudy (aaronandyson):

Yes.

OpenStudy (anonymous):

okay thank you

OpenStudy (aaronandyson):

Np!